A body of mass 2kg moving with velocity of 6m/s collides with a stationary object of mass 0.5kg. If the two bodies move together after the impact, calculate their common velocity.


1
Expert's answer
2021-11-05T18:20:01-0400

Given:

"m_1=2\:\rm kg", "m_2=0.5\:\rm kg"

"v_1=6\:{\rm m/s},\quad v_2=0"


The law of conservation of momentum gives


"m_1v_1+m_2v_2=(m_1+m_2)u""u=\frac{m_1v_1}{m_1+m_2}=\frac{2*6}{2+0.5}=4.8\;\rm m/s"